The demand function is given by D(p)= 5,000-100p, where p is the price and If the price is 20 birrs, calculate price elasticity of demand
1
Expert's answer
2020-10-07T07:15:32-0400

"D=5000-100p"

"q=5000-100p"

"dq\/dp=-100"



"p=20""q=5000-100\/20"

"q=3000"

"\in=|dq/dp/p/q|""\in=|-100 / 20/3000|"

"\in=2/3"

"\in=0.666"